### v1

# Fisher information, Crámer-Rao complexity and temperature in logistic map

Abstract
In this work we study the dynamics of the logistic map based on a probabilistic characterization in terms of the invariant density. We analyze the relevant regimes of the dynamics (regular, oscillatory, onset chaotic and fully chaotic) in terms of the Fisher information and the Crámer-Rao (CR) complexity. We found that these informational quantifiers allow to distinguish the dynamical regions of the map, by maximizing the Fisher information in the regular behavior and with the CR complexity exhibiting variations and a maximum near to the Pomeau-Maneville scenario. Fisher information as a function of time is examined to the light of the Frieden’s informational interpretation of the Thermodynamics Second Law. Finally, we apply the Equipartition Theorem to propose a definition of temperature for the logistic map, thus providing a macroscopic signature of the dynamics.
I Introduction
Maps constitute useful tools to study the discrete evolution of a system expressed by $x_{i+1}=f(x_{i})$, with $f(x)$ a function describing the dynamics. The logistic map constitutes one of the most interesting objects due to its several applications in physics, biology, chemistry, engineering science and mathematics [1]. The iteration of its quadratic formula reveals many types of dynamical regimes going from the regular ones to the fully chaotic one, and between them exhibiting intermittency, aperiodicity, and regular-chaotic mixed behaviors, thus manifesting the richness of the logistic map for testing chaos, ergodicity, onset of chaos, complexity and serving for population growth approaches, among others. Complementarily, a probabilistic description for discrete systems is the given by the theory of the invariant measures and Markov operators [2]. In this approach, the features of the dynamics are studied from the evolution of probabilistic densities that characterize the knowledge of the system at a given instant. The evolution of any density is obtained from the iterations of the Frobenius-Perron operator over an arbitrary initial density. Of particular importance is the existence of invariant densities of the Frobenius-Perron operator since they represent stationary states modeling the system in the asymptotic limit of large times.
On the other hand, statistical quantifiers have been shown to be important theoretical tools for characterizing several phenomena. Between these ones, the Fisher information (FI) has attracted special attention in the literature. FI arose from the problem of parameter estimation in the context of statistics, and since then it has been employed for characterizing pattern genetics in biology, dynamical phase transitions in environmental engineering, chemistry, economy, quantum mechanics, deduction of motion equations, by making use of its sensibility against local variations of the probability distribution. Intuitively, the Fisher information can distinguish situations with behaviors being radically different. For instance, a regular behavior characterized by a piked distribution gives a high valued Fisher information, while a fully chaotic dynamics represented by a uniform distribution minimizes the Fisher information. The CR complexity plays a complementary role by distinguishing behaviors where the Fisher information is not able for. When we have different dynamical regimes with similar values of the Fisher information, we can calculate the CR complexity for distinguishing between them. Additionally, FI has been studied to give an alternative formulation of the Second Law of Thermodynamics. Illustrating with several examples and theoretical arguments, in [3] Frieden has suggested that FI can provide a interpretation of the arrow of time, by postulating that Second Law of Thermodynamics is equivalent to the condition $dI/dt\leq 0$. Physically, this inequality means that if the FI level of a system is observed to be decreasing, that is $dI<0$, then its history is necessarily advancing, which is the main content of the “$I$-Theorem” (Section 1.8.2 of [3]).
The main goal of this work is to characterize some relevant regimes of the logistic map dynamics from a probabilistic description in terms of the invariant density, from which we calculate the Fisher information (FI) and its associated Cramér-Rao (CR) complexity for analyzing dynamical transitions. In addition, we propose a notion of temperature based on the Equipartition Theorem, which allows to provide a macroscopic signature of the map along with a study of the behavior of the Thermodynamics Second Law by means of the Frieden’s informational interpretation. The work is organized as follows.
II Preliminaries
We begin by presenting the minimal concepts and methods to be employed in the subsequent Sections of the paper.
II.1 Probability density functions, discrete maps and invariant density
A probability density function defined over an abstract space $\Gamma$ (typically, a subset of $\mathbb{R}^{m}$) is any nonnegative function $p:\Gamma\rightarrow\mathbb{R}_{+}$ such that $\int_{\Gamma}p(x)dx=1$ with $\Gamma$ called the space of events. In an experiment, $\Gamma$ represents the set of all possible outcomes. If these ones are $W$ possible results, then one has a discrete probability distribution expressed by a probability vector $(p_{1},p_{2},\ldots,p_{W})$ being $p_{i}$ the probability that the ith result occurs and the normalization condition reads $\sum_{i=1}^{W}p_{i}=1$.
Given a set $\Gamma$ and a continuous function $f:\Gamma\rightarrow\Gamma$, we say that the sequence $\{x_{n}\}_{n\in\mathbb{N}_{0}}\subseteq\Gamma$ such that
$x_{n+1}=f(x_{n})\quad\forall n\in\mathbb{N}_{0}$ (1)
defines a discrete map. Given $x_{0}\in\Gamma$ as the starting point, the sets $\{x_{n}:n\in\mathbb{N}_{0}\}$ are called orbits, where $x_{\infty}=\lim_{n\rightarrow\infty}x_{n}$ stands for the limit of a orbit. From the physical viewpoint, a discrete map models a system whose dynamics is given by iterating (1), where each iteration corresponds to a time step. If the system is initially in a state $x_{0}$, then $x_{n}$ represents the state after $n$ time steps. When there exists an element $c\in\Gamma$ such that $f(x^{*})=x^{*}$, we say that $x^{*}$ is a fixed point of $f$ and they are physically interpreted as stationary states of the system. The dynamics of a discrete map can be characterized in terms of probability density functions as follows. The Frobenius–Perron operator $P:\mathbb{L}^{1}(\Gamma)\rightarrow\mathbb{L}^{1}(\Gamma)$ associated with the map (1) is given by [2]
$\int_{A}P\phi(x)dx=\int_{f^{-1}(A)}\phi(x)dx$ (2)
for all $\phi\in\mathbb{L}^{1}(\Gamma)$ and $A\subseteq\Gamma$ with $f^{-1}(A)$ the preimage of $A$ and $\mathbb{L}^{1}(\Gamma)$ the set $\{g:\Gamma\rightarrow\mathbb{R}\ \textrm{with}\ \int_{\Gamma}|g(x)|dx<\infty\}$. Any non-negative normalized function $\rho\in\mathbb{L}^{1}(\Gamma)$ with
$P\rho(x)=\rho(x)\quad\forall x\in\Gamma$ (3)
is called the invariant density of the map, constituting a invariant measure. Numerically, the invariant density $\rho(x)$ can be constructed by dividing $\Gamma$ into $W$ intervals and defining
$\rho_{i}=\frac{\#\{x_{\tau}\in[\frac{i-1}{W},\frac{i}{W})\ |\ \tau=0,1,\ldots,N\}}{N}$ (4)
as the fraction of points $x_{\tau}$ in the interval $\frac{i-1}{W},\frac{i}{W})$ generated by iterations of $f(x)$ from an initial point $x_{0}\in\Gamma$. Thus, when $N\gg 1$ and $N\gg W$ the invariant density $\rho(x)$ results well approximated numerically by
$\rho(x)=\sum_{i=1}^{W}\rho_{i}1_{[\frac{i-1}{W},\frac{i}{W})}(x)$ (5)
with $1_{[\frac{i-1}{W},\frac{i}{W})}(x)$ standing for the characteristic function of the interval $\frac{i-1}{W},\frac{i}{W})$. It is worth to be noted that in the limit $N\rightarrow\infty$ the invariant density $\rho(x)$ does not depend on the starting point $x_{0}$.
II.2 Logistic map
The logistic map is given by [1]
$x_{n+1}=\mu x_{n}(1-x_{n})\quad,\quad n\in\mathbb{N}_{0}$ (6)
where $\mu\in(0,4]$ is the external parameter and $x_{0}\in\Gamma=[0,1]$ the initial condition. The case $\mu=0$ is not of interest since we have $x_{n+1}=0$ for all $n\in\mathbb{N}_{0}$. We review some relevant regimes that we will employ for characterizing them from the formalism presented. When $0<\mu\leq 3$ all the orbits are convergent and we have that $x_{\infty}=0$ for $0<\mu<1$ and $x_{\infty}=\frac{\mu-1}{\mu}$ for $1\leq\mu\leq 3$ with different rates of convergence. For $0<\mu<3$ the orbits rapidly approach to the asymptotic limit with a linear convergence for $2\leq\mu<3$ and less than linear for $\mu=3$. When $3\leq\mu\leq 3.44949$ the orbits oscillate between two values and with $3.44949<\mu\leq 3.56995$ almost all the orbits oscillate between four values. The value $\mu=3.56995$ represents the beginning of the chaotic behavior (onset of chaos) where orbit oscillations of finite period are not observed. The region $3.56995\leq\mu\leq 3.82843$ represents the called Pameau-Maneville scenario, with the orbits manifesting a periodic laminar phase interrupted by bursts of aperiodic behavior. For $\mu=4$ the fully chaotic behavior emerges with all the orbits dense in $[0,1]$ provided with a mixing dynamics. Interestingly, when $\mu=2$ and $\mu=4$ there exist exact solutions for the orbits, given by
$x_{n}=\begin{cases}\frac{1-\exp{(2^{n}\log(1-x_{0}))}}{2}&\textrm{for}\quad\mu=2\\ \frac{1-\cos{(2^{n}\arccos(1-x_{0}))}}{2}&\textrm{for}\quad\mu=4.\end{cases}$ (7)
The invariant density has a closed formula for $\mu=4$. Using (2), in this case we have for all $\phi\in\mathbb{L}^{1}([0,1])$
$P\phi(x)=\frac{\phi\Bigg(\frac{1}{2}+\frac{1}{2}\sqrt{1-x}\Bigg)+\phi\Bigg(\frac{1}{2}-\frac{1}{2}\sqrt{1-x}\Bigg)}{4\sqrt{1-x}},$ (8)
so in the limit of infinite successive iterations $N\rightarrow\infty$ we obtain the invariant density $\rho(x)$
$\rho(x)=\lim_{N\rightarrow\infty}P^{N}\phi(x)=\frac{1}{\pi\sqrt{x(1-x)}}.$ (9)
It should to be noted that, except for particular cases, the invariant density has no analytical closed formula so in general we have to compute it only numerically.
II.3 Fisher information and Crámer-Rao complexity
The discrete version of the Fisher (FI) information can be given in terms of the expression
$I[p]=4\sum_{i=0}^{W-1}\Big(\sqrt{p_{i+1}}-\sqrt{p_{i}}\Big)^{2}$ (10)
for all discrete probability density $\{p_{i}\}$, $p_{i}\geq 0$ and $\sum p_{i}=1$. The associated statistical quantifier to the Fisher information is the Crámer-Rao complexity, given by
$C[p]=I[p]\times\sigma^{2}$ (11)
where $\sigma^{2}$ stands for the variance of the probability density $\{p_{i}\}$, that is
$\sigma^{2}=\sum_{i=1}^{W}z_{i}^{2}p_{i}-\Big(\sum_{i=1}^{W}z_{i}p_{i}\Big)^{2}$ (12)
and $z_{i}\in[\frac{i-1}{W},\frac{i}{W})$ with $p=(p_{1},p_{2},\ldots,p_{W})$ a discrete probability density. The Crámer-Rao bound states that the CR complexity is low bounded by 1, i.e. $C[p]\geq 1$, being maximized for the Gaussian probability densities. We only focus on the Fisher information and the CR complexity so the employment of other measures of information and its associated complexities [4] is out of the scope of this work.
II.4 Equipartition Theorem
Equipartition Theorem (ET) establishes a relation between a temperature of a system and the average of the total kinetic energy of their particles, expressed in the general form by
$\Bigg\langle q_{m}\frac{\partial H}{\partial q_{n}}\Bigg\rangle=\delta_{mn}k_{B}T$ (13)
where the brackets means an ensemble average provided with the Liouville equilibrium density in the asymptotic limit $t\rightarrow\infty$, $H$ is the Hamiltonian of the system, $k_{B}$ is the Boltzmann constant, $T$ is the temperature and $q_{n}$ is the nth coordinate of the system in phase space. For the special case of a system composed by non-interacting particles of mass $m$, the average energy per particle $E_{\textrm{particle}}=\langle H_{\textrm{particle}}\rangle$ is given by the Hamiltonian $H_{particle}=(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})/2m$, so applying (13) for $q_{i}=p_{i}$ ($i=x,y,z$) it follows
$E_{\textrm{particle}}=\frac{m}{2}\langle v^{2}\rangle=\frac{1}{2}\sum_{i=x,y,z}\Bigg\langle p_{i}\frac{\partial H}{\partial p_{i}}\Bigg\rangle=\frac{3}{2}k_{B}T,$ (14)
from which we can case that temperature of the system is proportional to the average energy per particle. It is worth noting that some hypotheses are needed for the validity of ET. Typically, ET holds for ergodic systems in thermal equilibrium with a density distribution in phase space equiprobable for all the states. For our proposal it is convenient to make (14) dimensionless, so considering the one-dimensional case, from (14) we can deduce the dimensionless temperature $T/T_{0}$ as
$\frac{T}{T_{0}}=\frac{\langle v^{2}\rangle}{v_{0}^{2}}\frac{mv_{0}^{2}}{k_{B}T_{0}},$ (15)
where $T_{0}$ and $v_{0}$ are arbitrary characteristic temperature and velocity for the system.
III Statistical quantifiers and temperature for logistic map
We calculate the Fisher information and its CR complexity for a set of values of the parameter $\mu$ of the logistic map. Next, for some representative values of $\mu$, we analyze the temporal evolution of the Fisher information from the Frieden’s interpretation of Second Law of Thermodynamics. Finally, by employing the Equipartition Theorem an associated temperature for the logistic map is obtained along with the FI and the CR complexity as a function of the temperature.
III.1 Fisher information and CR complexity of logistic map
For the numerical calculations of the FI and the CR complexity we employed $N=10^{6}$ steps and $W=10^{4}$ bins, thus guaranteeing $N\gg W$. By increasing $N$ the results remained the same due to the convergence of the invariant density. In figures 1 and 2 the FI and the CR complexity in function of the parameter $\mu$ are illustrated.
Figure 1: Fisher information (10) of the invariant density (5) of the logistic map as a function of the parameter $\mu$ for $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{6}$ number of steps and $W=10^{4}$ bins.
Figure 2: Crámer-Rao complexity (11) of the invariant density (5) of the logistic map as a function of the parameter $\mu$ for $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{6}$ number of steps and $W=10^{4}$ bins.
From Fig. 2 we see that FI presents an absolute maximum value for $\mu=1$, which corresponds to regular behavior with orbits converging to $x_{\infty}=0$. For $0\leq\mu<1$ and $1\leq\mu<3.4$ the FI behaves approximately constant, from which it is not possible to distinguish the dynamical transitions for these values. From $\mu=3.4$ the FI decreases up to the region $3.8<\mu<3.9$ where a local maximum occurs. FI in the fully chaotic regime $\mu=4$ has the same value as for $3.7\leq\mu\leq 3.8$. By contrast, from Fig. 2 we see that the CR complexity provides a distinction between
In Fig. 3 we see how the CR complexity takes high values while the FI is vanishingly small, and vice-versa.
Figure 3: CR complexity (11) as a function of the Fisher information (10) for the invariant density (5) of the logistic map with $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{6}$ number of steps and $W=10^{4}$ bins.
III.2 Evolution of Fisher information and CR complexity of logistic map
With the aim of testing the Frieden’s informational interpretation of the Second Law we analyze the FI and the CR complexity as functions of time.
III.3 Temperature associated to logistic map
Considering that $v_{i}=x_{i+i}-x_{i}$ represents a dimensionless average velocity of the logistic map between the ith and (i+1)th steps, then we apply the relationship (15) with $mv_{0}^{2}/(k_{B}T_{0})=1$ and propose the map temperature (MP) at Nth step, denoted by $T(\mu,N,M)$ for the logistic map as follows:
$T(\mu,N,M)=\frac{1}{M}\sum_{j=1}^{M}(x_{N+1}^{(j)}-x_{N}^{(j)})^{2}\quad\forall\ N=0,1,\ldots,$ (16)
where $\{x_{k}^{(j)}\}$ is the orbit starting at $x_{0}^{(j)}$ with $j=1,\ldots,M$.
III.3.1 Numerical results
We calculate MP (16) for the same range of values of $\mu=0.05,0.1,\ldots,3.95,4$ as before, with $M=10^{5}$ initial conditions and $N=10^{3}$. It is worth to mention that for $M>10^{5}$ the MP remained with the same behavior. When calculating MP from definition (16), we observed the existence of a short time transient along with the presence of small variations of the $T(\mu,N,M)$ for some values of $\mu$. In order to define a temperature for a large number of steps we have taken an average neglecting the transient. The dependence of the MP in function of $\mu$ is shown in Fig. 4. For a better visualization, the MP was normalized according to $T(\mu,N,M)/\max\{T(\mu,N,M)\}$ with the the maximum taken for $N\geq N_{0}$, being $N\leq N_{0}$ the transient region.
Figure 4: Averaged value of the map temperature (16) for $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{3}$ number of steps and $M=10^{5}$ initial conditions.
III.3.2 Temperature as a function of time
III.3.3 Some formal results
We can demonstrate the following formal result for the map temperature MP (16).
Theorem: For the function $T(\mu,N)$ we have:
  • $(I)$
    If $0<\mu\leq 3$ then $T(\mu,N)\rightarrow 0$ when $N\rightarrow\infty$.
  • $(II)$
    If $\mu=4$ then $|T(4,N)-K|<\varepsilon$ for all $N\geq C$ with $C>0$.
Proof: In order to bound $\cos(2^{n+1}\arccos{(1-x_{0}^{(j)})})-\cos(2^{n}\arccos{(1-x_{0}^{(j)})})]^{2}$ we note that this expression can be recasted as $(\cos(2(2^{n}y_{j}))-\cos(2^{n}y_{j}))^{2}$ with $y_{j}=\arccos{(1-x_{0}^{(j)})}$. In turn, we have $(\cos(2(2^{n}y_{j}))-\cos(2^{n}y_{j}))^{2}=(1/4)(e^{i(2(2^{n}y_{j}))}-e^{-i(2(2^{n}y_{j}))}-(e^{i(2^{n}y_{j})}-e^{i(2^{n}y_{j})}))^{2}$
III.4 Fisher information and CR complexity as a function of temperature
In […] it has been shown that the Fisher information of the Liouville distribution $\rho(q,p)$ of the canonical ensemble is a decreasing monotonic function of temperature. Inspired by this feature, now we study the FI and the CR as functions of the temperature. In Fig.
III.5 Fisher information and CR complexity as a function of time
IV Conclusions
The results of this work are potentially extended to other chaotic maps, for instance the tent map and the circular map.
Acknowledgments
Samuel S. Santos, Guilherme V. B. Junior and Ignacio S. Gomez acknowledge support from the Department of Exact and Natural Sciences of the State University of Southwest Bahia (UESB), Itapetinga, Bahia, Brazil. Ignacio S. Gomez acknowledges support received from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant Number 316131/2023-7. Guilherme Vieira Brito Junior acknowledges support received from FAPESB as a scientific initiation scholarship.